Circular No. 6336

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) C. M. Lisse and M. F. A’Hearn, University of Maryland, report obser-

vations of comet C/1996 B2 from the NASA Infrared Telescope Facility on Feb. 7.5 UT (r = 1.90 AU, Δ = 1.57 AU): “Using a 5′′.6 aperture, total magnitudes are J = 12.0 ± 0.02, H = 11.6 ± 0.02, K = 11.5 ± 0.02 (2σ errors) — slightly redder than the sun, suggesting the lack of icy grains — and N = 4.06 ± 0.16 (2σ). The J − N color is unusually small, suggesting either a very low thermal emissivity or a very high albedo for the dust.”

4U 1608–52 J. van Paradijs, University of Alabama in Huntsville and University of

Amsterdam; W. Zhang, Universities Space Research Association (USRA) and Goddard Space Flight Center (GSFC); F. Marshall and J. H. Swank, GSFC; T. Augusteijn, European Southern Observatory; E. Kuulkers, European Space Agency; W. H. G. Lewin, Massachusetts Institute of Technology; F. Lamb, University of Illinois; I. Lapidus, University of Cambridge; J. Lochner and T. Strohmayer, USRA/GSFC; M. van der Klis, University of Amsterdam; B. Vaughan, California Institute of Technology; report: “Following the recent outburst of 4U 1608–52 (IAUC 6331), we have observed this low-mass x-ray binary for 3 hr on Mar. 3 with the Rossi XTE (total exposure ∼ 2 hr). The source showed strong quasi-periodic intensity oscillations (QPOs) throughout the observation. In the beginning, the QPOs were observed with a centroid frequency of 850 Hz, a FWHM of 6.8 Hz, and a fractional rms of 6.5 percent. Twenty minutes into the observation, the centroid frequency started increasing and 15 min later reached 890 Hz, with a fractional rms value of 5.2 percent and a FWHM of 15 Hz. Then the centroid frequency started decreasing through the end of the observation, when it reached 828 Hz with a fractional RMS of 6.8 percent and a FWHM of 6.7 Hz. In the course of the observation, the source intensity in the XTE/PCA band (2–60 keV) showed a slight overall decrease from 3200 counts/s in the beginning to 3100 counts/s in the end. The excursion to higher frequency did not correlate with any change in intensity. With earlier detections of kHz QPOs in the source 4U 1728–34 (IAUC 6320) and the Z source Sco X-1 (IAUC 6319), our results suggest that kHz QPO are a regular feature of low-mass x-ray binaries.”

1996 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green